# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v2

@@ A.1 Remarks on Shape Diameter Function (ShDF)
Our semantic-aware parameterization, explained in Sec.3.3, relies on computing a per-vertex semantic partition of the input mesh using shape diameter function Shapira et al. (2008). To estimate per-vertex semantic label, we design and implement a method based on Shape Diameter Function Shapira et al. (2008) that involves the following key steps.
Compute a per face local thickness (shape-diameter) field. For a surface sample $p$ (face centroid or vertex), the value $\mathrm{ShDF}(p)$ estimates the local object diameter by probing the interior along a cone centered on the inward normal at $p$. Let $\mathcal{R}(p)=\{r_{i}\}_{i=1}^{R}$ be a set of rays radiating from $p$ inside a cone of half-angle $\alpha$ around the inward normal. For each ray $r_{i}$ we compute the length $\ell_{i}$ of the intersection segment between $r_{i}$ and the mesh interior (i.e., the distance from $p$ to the first opposite intersection). We robustly aggregate the ensemble $\{\ell_{i}\}$ using a median (or trimmed mean) to obtain: $\mathrm{ShDF}(p)\;=\;\mathrm{median}\big(\{\ell_{i}\}_{i=1}^{R}\big).$ This median-based aggregation reduces sensitivity to spurious long or short intersections caused by complex local geometry. In practice, to estimate a scalar per surface element that measures local object thickness (intuitively, it presents the diameter of the object in the neighborhood of each sample), we use ray casting (using PyMeshLab Muntoni & Cignoni (2021)) in multiple directions per sample and return a per-face scalar. Then, the raw scalars are normalized to $[0,1]$ and re-scaled via a monotonic compression (e.g., a log-like transform) to reduce dynamic range and improve cluster separability. In our implementation we use per-face ShDF evaluated at face centroids with the following defaults: cone angle $\theta=120^{\circ}$ (as in the original work Shapira et al. (2008)), number of rays $R=60$, and ray sampling distributed uniformly within the cone.
@@
Similar to Shapira et al. (2008), we smooth raw ShDF samples on the mesh to promote spatial coherence and reduce pose-dependent noise. We apply a small number (e.g., 1-2) of iterations of Laplacian smoothing on the ShDF scalar field with area-weighted averaging. To balance values across scales and to boost contrast of small parts, we perform a normalization:
@@ A.2 Remarks on Base Architecture for UV-Parameterization Learning
Our backbone, based on bi-directional cycle mapping proposed in Zhao et al. (2025), implements a learnable and interpretable pipeline that mimics the artist operations used to create a UV atlas: (i) deforming network, starting from a regular 2D grid of uniformly spaced coordinates, this network adaptively deforms the grid to produce a set of candidate UV coordinates that are potentially well-suited for mapping; (ii) wrapping network, this module maps 2D UV candidates, generated by the deforming network, to the 3D surface by smoothly bending and folding the deformed planar grid until it conforms to the target geometry (in this case the input 3D mesh) which leads to effectively bridge the 2D and 3D domains; (iii) cutting network, operating directly on the 3D surface, generated by the wrapping network, this network transforms the input closed 3D geometry, obtained from the wrapping network, into an open, more developable 3D surface manifold. By strategically determining seam locations, i.e. where to cut the input closed 3D surface, it facilitates distortion reduction in subsequent flattening steps; and (iv) unwrapping network, this final network projects the open 3D surface points back to the 2D domain, preserving smoothness and geometric coherence. In this design, each subnetwork is designed as a multilayer perceptron (MLP) applied pointwise to mesh vertices. The entire architecture is trained jointly in an end-to-end unsupervised fashion. The four subnetworks are assembled into two complementary branches that enforce both 2D-3D-2D and 3D-2D-3D cycle mappings, ensuring consistency across 2D and 3D domains. Coupled with differentiable loss functions designed to encourage bijectivity and minimize distortion, this backbone forms the base architecture of our method to achieve a candidate UV parameterization for the next stages. In the following, we detail on the loss designs and the interaction of the four sub-networks within the two-branch bi-directional cycle.
@@
Bi-Directional Cycle Mapping. We follow Zhang et al. (2024b); Srinivasan et al. (2024); Zhao et al. (2025) to build our base neural UV parameterization architecture with two-branch MLP architecture that leverages the four sub-networks, explained above.
@@
The DDL is computed by applying automatic differentiation to the MLP outputs to approximate the local Jacobian, while the TDL explicitly measures changes in angles and areas between 3D triangles and their corresponding 2D counterparts. In practice, we adopt the same formulations of DDL and TDL as proposed in Zhao et al. (2025), and refer readers to Sheffer et al. (2007b); Hormann et al. (2007); Floater & Hormann (2005) for further details on distortion metrics and differential geometry concepts.
@@ A.3 Additional Experiments and Results
Ablation studies on different UV packing (aggregator) methods. To further improve our semantic-aware UV parameterization pipeline’s output, a simple enhancement is to post-process the UV charts by jointly optimizing two key UV metrics: texel density and UV space utilization. Higher texel density improves the quality of the projected 3D texture, and maintaining consistent texel density across charts ensures uniform visual fidelity. At the same time, efficient space utilization maximizes the use of available texture memory. Hence, to analyze the effect of more advanced UV-chart packing (aggregator) strategies (different from the one proposed in Sec. 3.3), we perform an ablation study and evaluate several UV-packing techniques from three widely used commercial tools including, Autodesk Maya Autodesk, Inc. (2025), Houdini Toronto-based SideFX (2025), and Blender Blender Foundation (2025). We use a 3D cow mesh shown in Fig. 7 to evaluate texel density and UV-space utilization of UV parameterizations packed using different UV packing strategies from the three commercial tools. Specifically, we apply our semantic-aware UV parameterization framework as follows: (i) the 3D cow is first partitioned by PartField Liu et al. (2025) into seven semantic 3D parts, (ii) then the geometry-preserving neural backbone is applied separately to each part to learn UV parameterizations for each 3D part, (iii) instead of using our default UV packing (presented in Sec. 3.3), we apply different UV packing strategies from the three commercial tools, and (iv) lastly, for each packing strategy, we compute two UV metrics, namely texel density and UV-space utilization, for the UV parameterizations resulting from the three different UV packing methods. Table 5 reports the resulting values for both metrics to illustrate how each tool performs on a 3D cow mesh example. Also, in Fig. 7, we qualitatively compare the three commercial tools, applying their automatic packing and layout features to the same unwrapped mesh of 3D cow.
@@
Method Texel Density UV Space Utilization Chart1 Chart2 Chart3 Chart4 Chart5 Chart6 Chart7 Avg Maya Autodesk, Inc. (2025) 160.62 150.41 150.41 150.41 156.51 150.41 150.41 152.74 71% Blender Blender Foundation (2025) 151.66 142.02 142.02 142.02 147.78 142.02 142.02 144.22 57% Houdini Toronto-based SideFX (2025) 157.08 147.09 147.09 147.09 153.06 147.09 147.10 149.37 62%
@@
Figure 7: Qualitative ablations of different UV-packing strategies. To further improve our semantic-aware UV parameterization pipeline’s output, we present alternative more advanced UV packing results (different from the one proposed in Sec. 3.3) obtained from three widely used commercial tools: (a) Autodesk Maya, (b) Blender, and (c) Houdini. The left-most column shows the 3D cow mesh segmented by PartField Liu et al. (2025). In the Autodesk Maya result (a), each UV chart corresponding to a semantic 3D part is colored to match its counterpart on the 3D mesh. The same applies to the Blender and Houdini results. Specifically, the 3D cow is first partitioned by PartField Liu et al. (2025) into seven semantic 3D parts. Then, the geometry-preserving neural backbone (Sec. 3.2) is applied separately to each part to learn UV parameterizations for each 3D part leading to seven UV charts. Then, instead of using our default UV packing (presented in Sec. 3.3), we apply automatic packing strategies layout features of the three commercial tools to the unwrapped mesh of 3D cow.
@@
Ablation studies on different partitioning strategies (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize (semantic-aware UV parameterization) pipeline identical. Then, we train each pipeline separately on input 3D meshes and measure and report angle-preservation (conformality) and area-preservation (equiareality) properties in the generated UV parameterization. Figs. 1415 and Table 6 shows the qualitative and quantitative results of these experiments, respectively. To interpret the numbers in the table, we point out two points: (i) There is no obvious or general relationship between the choice of segmentation algorithm and the geometry-preservation metrics (conformality and equiareality) of the final UV parameterization. This is due to the fact that existing 3D segmentation methods were developed to produce semantically meaningful parts, not to optimize downstream UV parameterization. Consequently, they do not incorporate UV-specific information. Therefore, from the numbers in Table 6, we cannot draw a unified, causal conclusion connecting segmentation method type to final UV geometry properties. Having that said, we include this segmentation ablation and Table 6 intentionally to inform practitioners that different partitioning methods, while varying in semantic quality, can also influence geometry-preservation in the resulting UVs. Reporting these results shows the important trade-off between semantic-awareness and geometric fidelity. Therefore, practitioners should try different segmentation approaches and choose the one that best fits their application needs. (ii) Of course, state-of-the-art, learning-based segmentation methods can produce more semantically coherent partitions than ShDF Shapira et al. (2008) (as shown by works such as SAMesh Tang et al. (2024)). Therefore, any improvements in that literature can directly improve the semantic-awareness of our pipeline because segmentation is the first stage of our semantic-aware parameterization (Fig.2). In other words, a better 3D partitioner is an immediate plug-in improvement for our method.
@@
Analysis on training and inference computation times. To provide users and practitioners with some insights about the computation time needed for the proposed method in this study, we report both inference and training time of the proposed method in Table 3, Table 1, Table 8, and Table 9, and compare them with the computation time of the state-of-the-art method (at the time of writing this manuscript), FlexPara Zhao et al. (2025).
@@
  • Inference time. (i) Visibility-aware pipeline: although training is performed per-mesh (as is common in this literature: Nuvo Srinivasan et al. (2024), FlexPara Zhao et al. (2025), OptCuts Li et al. (2018)), the visibility objective is applied only during training (not the inference stage). Specifically, at inference the trained model uses the same neural architecture as FlexPara Zhao et al. (2025) and therefore incurs no additional runtime overhead. This is why the inference times for our Visibility-Aware Param and FlexPara are the same in Table 1 ($\approx$2 seconds). (ii) Semantic-aware pipeline: the semantic-aware pipeline does introduce extra inference cost because it performs three stages at test time: (a) 3D semantic partitioning, (b) per-part parameterization (running the backbone once per part), and (c) aggregation/UV packing. These additional steps (segmentation, multiple backbone invocations, and packing) explain the larger inference time reported in Table 2 and for the combined pipeline (semantic+visibility-aware) in Table 1. A simple and effective mitigation is to parallelize the per-part backbone calls by using multi-threading during the inference, which would significantly reduce inference time.
@@
  • Training time. (i) Semantic-aware pipeline (Table 8): counterintuitively, our semantic-aware pipeline achieves shorter training time than FlexPara Zhao et al. (2025) across different 3D meshes. The reason is that the mesh is partitioned into many smaller submeshes: although our semantic-aware pipeline calls the backbone multiple times (once per semantic part), which might suggest longer training, this is compensated by the fact that the 3D segmented parts are much smaller than the original input mesh used by FlexPara. Hence, training on many small submeshes is faster than training a single model on a large, complex mesh with many vertices and faces. This empirical observation is consistent with our careful time measurement of each stage in the semantic-aware pipeline which is: learning UV parameterization for small 3D parts is much faster than doing the same for large meshes. More importantly, the reduction in per-part training cost outweighs the multiplicative overhead of multiple backbone runs, which results in shorter overall training time for the semantic-aware pipeline in our experiments. This is why Table 8 shows shorter training times for our semantic-aware pipeline compared to FlexPara in our experiments. (ii) Visibility-aware pipeline (Table 9): The visibility-aware pipeline trains slower than FlexPara Zhao et al. (2025) across different meshes. This is expected because, compared to the base neural backbone (i.e., FlexPara), our visibility-aware pipeline introduces per-iteration overheads: (a) per-vertex ambient occlusion (AO) computation, (b) candidate UV generation via the backbone, (c) seam/boundary extraction in UV space, and (d) AO-weighted loss computation. The most expensive stage is seam extraction which accounts for approximately 90% of the additional training time. Specifically, for each vertex we inspect its 3–5 neighbors, compare mapped UV positions, and evaluate Eq. 5 to detect UV discontinuities. This neighbor-checking and distance computation scales with vertex count and dominates the additional training cost ( 90% of the overhead), making visibility-aware training significantly heavier for complex meshes.
@@
Table 6: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Quantitatively, using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh.
@@
Table 8: Comparison of training time between our proposed semantic-aware parameterization pipeline and FlexPara’s multi-chart parameterization Zhao et al. (2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Rabbit Potion Squidward House Average FlexPara (multi-Chart) Zhao et al. (2025) 900 (0.0900) 16377 (1.6377) 13784 (1.3784) 10353 (1.0353) Ours - Semantic-Aware Param 676 (0.1352) 666 (0.1332) 602 (0.1205) 648 (0.1296)
Table 9: Comparison of training time between our proposed visibility-aware parameterization pipeline and FlexPara’s global (single-chart) parameterization Zhao et al. (2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Duckie Fish FlipFlop Rabbit Spot Average FlexPara (single-Chart) Zhao et al. (2025) 1306 (0.1306) 1323 (0.1323) 1275 (0.1275) 1293 (0.1293) 1230 (0.1230) 1285 (0.1285) Ours - Visibility-Aware Param 36642 (3.6642) 47460 (4.7360) 16864 (3.3729) 35232 (2.3488) 19605 (1.9605) 31160 (3.1160)
@@
Method General User Preference Percentage Visibility-Awareness Evaluation OptCuts Li et al. (2018) 1.43 % FlexPara (single-Chart) Zhao et al. (2025) 7.14 % Ours - Visibility-Aware Param 91.42 % Semantic-Awareness Evaluation xatlas jpcy (2025) 4.28 % Blender Blender Foundation (2025) 4.28 % Autodesk Maya Autodesk, Inc. (2025) 10 % FlexPara (multi-Chart) Zhao et al. (2025) 7.14 % Ours - Semantic-Aware Param 74.29 %
Figure 8: Quantitative distribution of seam-vertex ambient occlusion (AO) for three meshes parameterized with our method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018). Red bars indicate AO distributions of seam vertices, while black bars represent all vertices. The $y$-axis denotes vertex density and the $x$-axis AO values. Our method shifts seam-vertex AO distributions toward lower values (left side), showing that the visibility-aware pipeline relocates seams to less visible (more occluded) regions, unlike FlexPara and OptCuts.
Figure 9: Extended version of Fig. 5 in the main paper. Checkerboard texturing comparison using UV parameterizations produced by our visibility-aware method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018). Each row shows rendered views of different meshes textured with a checkerboard and a magnified inset of a visually important region near seams (red circles). Because our method steers seams toward occluded regions, the checkerboard pattern appears substantially more continuous from typical camera viewpoints. By contrast, baselines exhibit visible seam artifacts in the zoomed-in insets.
Figure 10: Extended version of Fig. 4 in the main paper. Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018) (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.
Figure 11: Extended version of Fig. 4 in the main paper. Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018) (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.

### v3

@@ 5 Conclusions
Acknowledgments
This research was supported by Autodesk Research and the Autodesk AI Lab. We are especially thankful for the collaborative environment, technical resources, and insightful discussions that contributed to shaping this work.
@@ A.1 Remarks on Shape Diameter Function (ShDF)
Our semantic-aware parameterization, explained in Sec.3.3, relies on computing a per-vertex semantic partition of the input mesh using shape diameter function (Shapira et al., 2008). To estimate per-vertex semantic label, we design and implement a method based on Shape Diameter Function (Shapira et al., 2008) that involves the following key steps.
Compute a per face local thickness (shape-diameter) field. For a surface sample $p$ (face centroid or vertex), the value $\mathrm{ShDF}(p)$ estimates the local object diameter by probing the interior along a cone centered on the inward normal at $p$. Let $\mathcal{R}(p)=\{r_{i}\}_{i=1}^{R}$ be a set of rays radiating from $p$ inside a cone of half-angle $\alpha$ around the inward normal. For each ray $r_{i}$ we compute the length $\ell_{i}$ of the intersection segment between $r_{i}$ and the mesh interior (i.e., the distance from $p$ to the first opposite intersection). We robustly aggregate the ensemble $\{\ell_{i}\}$ using a median (or trimmed mean) to obtain: $\mathrm{ShDF}(p)\;=\;\mathrm{median}\big(\{\ell_{i}\}_{i=1}^{R}\big).$ This median-based aggregation reduces sensitivity to spurious long or short intersections caused by complex local geometry. In practice, to estimate a scalar per surface element that measures local object thickness (intuitively, it presents the diameter of the object in the neighborhood of each sample), we use ray casting (using PyMeshLab (Muntoni & Cignoni, 2021)) in multiple directions per sample and return a per-face scalar. Then, the raw scalars are normalized to $[0,1]$ and re-scaled via a monotonic compression (e.g., a log-like transform) to reduce dynamic range and improve cluster separability. In our implementation we use per-face ShDF evaluated at face centroids with the following defaults: cone angle $\theta=120^{\circ}$ (as in the original work (Shapira et al., 2008)), number of rays $R=60$, and ray sampling distributed uniformly within the cone.
@@
Similar to (Shapira et al., 2008), we smooth raw ShDF samples on the mesh to promote spatial coherence and reduce pose-dependent noise. We apply a small number (e.g., 1-2) of iterations of Laplacian smoothing on the ShDF scalar field with area-weighted averaging. To balance values across scales and to boost contrast of small parts, we perform a normalization:
@@ A.2 Remarks on Base Architecture for UV-Parameterization Learning
Our backbone, based on bi-directional cycle mapping proposed in (Zhao et al., 2025), implements a learnable and interpretable pipeline that mimics the artist operations used to create a UV atlas: (i) deforming network, starting from a regular 2D grid of uniformly spaced coordinates, this network adaptively deforms the grid to produce a set of candidate UV coordinates that are potentially well-suited for mapping; (ii) wrapping network, this module maps 2D UV candidates, generated by the deforming network, to the 3D surface by smoothly bending and folding the deformed planar grid until it conforms to the target geometry (in this case the input 3D mesh) which leads to effectively bridge the 2D and 3D domains; (iii) cutting network, operating directly on the 3D surface, generated by the wrapping network, this network transforms the input closed 3D geometry, obtained from the wrapping network, into an open, more developable 3D surface manifold. By strategically determining seam locations, i.e. where to cut the input closed 3D surface, it facilitates distortion reduction in subsequent flattening steps; and (iv) unwrapping network, this final network projects the open 3D surface points back to the 2D domain, preserving smoothness and geometric coherence. In this design, each subnetwork is designed as a multilayer perceptron (MLP) applied pointwise to mesh vertices. The entire architecture is trained jointly in an end-to-end unsupervised fashion. The four subnetworks are assembled into two complementary branches that enforce both 2D-3D-2D and 3D-2D-3D cycle mappings, ensuring consistency across 2D and 3D domains. Coupled with differentiable loss functions designed to encourage bijectivity and minimize distortion, this backbone forms the base architecture of our method to achieve a candidate UV parameterization for the next stages. In the following, we detail on the loss designs and the interaction of the four sub-networks within the two-branch bi-directional cycle.
@@
Bi-Directional Cycle Mapping. We follow (Zhang et al., 2024b; Srinivasan et al., 2024; Zhao et al., 2025) to build our base neural UV parameterization architecture with two-branch MLP architecture that leverages the four sub-networks, explained above.
@@
The DDL is computed by applying automatic differentiation to the MLP outputs to approximate the local Jacobian, while the TDL explicitly measures changes in angles and areas between 3D triangles and their corresponding 2D counterparts. In practice, we adopt the same formulations of DDL and TDL as proposed in (Zhao et al., 2025), and refer readers to (Sheffer et al., 2007b; Hormann et al., 2007; Floater & Hormann, 2005) for further details on distortion metrics and differential geometry concepts.
@@ A.3 Additional Experiments and Results
Ablation studies on different UV packing (aggregator) methods. To further improve our semantic-aware UV parameterization pipeline’s output, a simple enhancement is to post-process the UV charts by jointly optimizing two key UV metrics: texel density and UV space utilization. Higher texel density improves the quality of the projected 3D texture, and maintaining consistent texel density across charts ensures uniform visual fidelity. At the same time, efficient space utilization maximizes the use of available texture memory. Hence, to analyze the effect of more advanced UV-chart packing (aggregator) strategies (different from the one proposed in Sec. 3.3), we perform an ablation study and evaluate several UV-packing techniques from three widely used commercial tools including, Autodesk Maya (Autodesk, Inc., 2025), Houdini (Toronto-based SideFX, 2025), and Blender (Blender Foundation, 2025). We use a 3D cow mesh shown in Fig. 6 to evaluate texel density and UV-space utilization of UV parameterizations packed using different UV packing strategies from the three commercial tools. Specifically, we apply our semantic-aware UV parameterization framework as follows: (i) the 3D cow is first partitioned by PartField (Liu et al., 2025) into seven semantic 3D parts, (ii) then the geometry-preserving neural backbone is applied separately to each part to learn UV parameterizations for each 3D part, (iii) instead of using our default UV packing (presented in Sec. 3.3), we apply different UV packing strategies from the three commercial tools, and (iv) lastly, for each packing strategy, we compute two UV metrics, namely texel density and UV-space utilization, for the UV parameterizations resulting from the three different UV packing methods. Table 5 reports the resulting values for both metrics to illustrate how each tool performs on a 3D cow mesh example. Also, in Fig. 6, we qualitatively compare the three commercial tools, applying their automatic packing and layout features to the same unwrapped mesh of 3D cow.
@@
Method Texel Density UV Space Utilization Chart1 Chart2 Chart3 Chart4 Chart5 Chart6 Chart7 Avg Maya (Autodesk, Inc., 2025) 160.62 150.41 150.41 150.41 156.51 150.41 150.41 152.74 71% Blender (Blender Foundation, 2025) 151.66 142.02 142.02 142.02 147.78 142.02 142.02 144.22 57% Houdini (Toronto-based SideFX, 2025) 157.08 147.09 147.09 147.09 153.06 147.09 147.10 149.37 62%
@@
Figure 6: Qualitative ablations of different UV-packing strategies. To further improve our semantic-aware UV parameterization pipeline’s output, we present alternative more advanced UV packing results (different from the one proposed in Sec. 3.3) obtained from three widely used commercial tools: (a) Autodesk Maya, (b) Blender, and (c) Houdini. The left-most column shows the 3D cow mesh segmented by PartField (Liu et al., 2025). In the Autodesk Maya result (a), each UV chart corresponding to a semantic 3D part is colored to match its counterpart on the 3D mesh. The same applies to the Blender and Houdini results. Specifically, the 3D cow is first partitioned by PartField (Liu et al., 2025) into seven semantic 3D parts. Then, the geometry-preserving neural backbone (Sec. 3.2) is applied separately to each part to learn UV parameterizations for each 3D part leading to seven UV charts. Then, instead of using our default UV packing (presented in Sec. 3.3), we apply automatic packing strategies layout features of the three commercial tools to the unwrapped mesh of 3D cow.
@@
Ablation studies on different partitioning strategies (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner (Shapira et al., 2008) with SAMesh (Tang et al., 2024) and PartField (Liu et al., 2025) while keeping the remainder of the partition-and-parameterize (semantic-aware UV parameterization) pipeline identical. Then, we train each pipeline separately on input 3D meshes and measure and report angle-preservation (conformality) and area-preservation (equiareality) properties in the generated UV parameterization. Figs. 1314 and Table 6 shows the qualitative and quantitative results of these experiments, respectively. To interpret the numbers in the table, we point out two points: (i) There is no obvious or general relationship between the choice of segmentation algorithm and the geometry-preservation metrics (conformality and equiareality) of the final UV parameterization. This is due to the fact that existing 3D segmentation methods were developed to produce semantically meaningful parts, not to optimize downstream UV parameterization. Consequently, they do not incorporate UV-specific information. Therefore, from the numbers in Table 6, we cannot draw a unified, causal conclusion connecting segmentation method type to final UV geometry properties. Having that said, we include this segmentation ablation and Table 6 intentionally to inform practitioners that different partitioning methods, while varying in semantic quality, can also influence geometry-preservation in the resulting UVs. Reporting these results shows the important trade-off between semantic-awareness and geometric fidelity. Therefore, practitioners should try different segmentation approaches and choose the one that best fits their application needs. (ii) Of course, state-of-the-art, learning-based segmentation methods can produce more semantically coherent partitions than ShDF (Shapira et al., 2008) (as shown by works such as SAMesh (Tang et al., 2024)). Therefore, any improvements in that literature can directly improve the semantic-awareness of our pipeline because segmentation is the first stage of our semantic-aware parameterization (Fig.2). In other words, a better 3D partitioner is an immediate plug-in improvement for our method.
@@
Analysis on training and inference computation times. To provide users and practitioners with some insights about the computation time needed for the proposed method in this study, we report both inference and training time of the proposed method in Table 3, Table 1, Table 8, and Table 9, and compare them with the computation time of the state-of-the-art method (at the time of writing this manuscript), FlexPara (Zhao et al., 2025).
@@
  • Inference time. (i) Visibility-aware pipeline: although training is performed per-mesh (as is common in this literature: Nuvo (Srinivasan et al., 2024), FlexPara (Zhao et al., 2025), OptCuts (Li et al., 2018)), the visibility objective is applied only during training (not the inference stage). Specifically, at inference the trained model uses the same neural architecture as FlexPara (Zhao et al., 2025) and therefore incurs no additional runtime overhead. This is why the inference times for our Visibility-Aware Param and FlexPara are the same in Table 1 ($\approx$2 seconds). (ii) Semantic-aware pipeline: the semantic-aware pipeline does introduce extra inference cost because it performs three stages at test time: (a) 3D semantic partitioning, (b) per-part parameterization (running the backbone once per part), and (c) aggregation/UV packing. These additional steps (segmentation, multiple backbone invocations, and packing) explain the larger inference time reported in Table 2 and for the combined pipeline (semantic+visibility-aware) in Table 1. A simple and effective mitigation is to parallelize the per-part backbone calls by using multi-threading during the inference, which would significantly reduce inference time.
@@
  • Training time. (i) Semantic-aware pipeline (Table 8): counterintuitively, our semantic-aware pipeline achieves shorter training time than FlexPara (Zhao et al., 2025) across different 3D meshes. The reason is that the mesh is partitioned into many smaller submeshes: although our semantic-aware pipeline calls the backbone multiple times (once per semantic part), which might suggest longer training, this is compensated by the fact that the 3D segmented parts are much smaller than the original input mesh used by FlexPara. Hence, training on many small submeshes is faster than training a single model on a large, complex mesh with many vertices and faces. This empirical observation is consistent with our careful time measurement of each stage in the semantic-aware pipeline which is: learning UV parameterization for small 3D parts is much faster than doing the same for large meshes. More importantly, the reduction in per-part training cost outweighs the multiplicative overhead of multiple backbone runs, which results in shorter overall training time for the semantic-aware pipeline in our experiments. This is why Table 8 shows shorter training times for our semantic-aware pipeline compared to FlexPara in our experiments. (ii) Visibility-aware pipeline (Table 9): The visibility-aware pipeline trains slower than FlexPara (Zhao et al., 2025) across different meshes. This is expected because, compared to the base neural backbone (i.e., FlexPara), our visibility-aware pipeline introduces per-iteration overheads: (a) per-vertex ambient occlusion (AO) computation, (b) candidate UV generation via the backbone, (c) seam/boundary extraction in UV space, and (d) AO-weighted loss computation. The most expensive stage is seam extraction which accounts for approximately 90% of the additional training time. Specifically, for each vertex we inspect its 3–5 neighbors, compare mapped UV positions, and evaluate Eq. 5 to detect UV discontinuities. This neighbor-checking and distance computation scales with vertex count and dominates the additional training cost ( 90% of the overhead), making visibility-aware training significantly heavier for complex meshes.
@@
Table 6: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField). We replace our default Shape Diameter Function (ShDF) partitioner (Shapira et al., 2008) with SAMesh (Tang et al., 2024) and PartField (Liu et al., 2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Quantitatively, using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh.
@@
Table 8: Comparison of training time between our proposed semantic-aware parameterization pipeline and FlexPara’s multi-chart parameterization (Zhao et al., 2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Rabbit Potion Squidward House Average FlexPara (multi-Chart) (Zhao et al., 2025) 900 (0.0900) 16377 (1.6377) 13784 (1.3784) 10353 (1.0353) Ours - Semantic-Aware Param 676 (0.1352) 666 (0.1332) 602 (0.1205) 648 (0.1296)
Table 9: Comparison of training time between our proposed visibility-aware parameterization pipeline and FlexPara’s global (single-chart) parameterization (Zhao et al., 2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Duckie Fish FlipFlop Rabbit Spot Average FlexPara (single-Chart) (Zhao et al., 2025) 1306 (0.1306) 1323 (0.1323) 1275 (0.1275) 1293 (0.1293) 1230 (0.1230) 1285 (0.1285) Ours - Visibility-Aware Param 36642 (3.6642) 47460 (4.7360) 16864 (3.3729) 35232 (2.3488) 19605 (1.9605) 31160 (3.1160)
@@
Method General User Preference Percentage Visibility-Awareness Evaluation OptCuts (Li et al., 2018) 1.43 % FlexPara (single-Chart) (Zhao et al., 2025) 7.14 % Ours - Visibility-Aware Param 91.42 % Semantic-Awareness Evaluation xatlas (jpcy, 2025) 4.28 % Blender (Blender Foundation, 2025) 4.28 % Autodesk Maya (Autodesk, Inc., 2025) 10 % FlexPara (multi-Chart) (Zhao et al., 2025) 7.14 % Ours - Semantic-Aware Param 74.29 %
Figure 7: Quantitative distribution of seam-vertex ambient occlusion (AO) for three meshes parameterized with our method, FlexPara (Zhao et al., 2025), and OptCuts (Li et al., 2018). Red bars indicate AO distributions of seam vertices, while black bars represent all vertices. The $y$-axis denotes vertex density and the $x$-axis AO values. Our method shifts seam-vertex AO distributions toward lower values (left side), showing that the visibility-aware pipeline relocates seams to less visible (more occluded) regions, unlike FlexPara and OptCuts.
Figure 8: Extended version of Fig. 5 in the main paper. Checkerboard texturing comparison using UV parameterizations produced by our visibility-aware method, FlexPara (Zhao et al., 2025), and OptCuts (Li et al., 2018). Each row shows rendered views of different meshes textured with a checkerboard and a magnified inset of a visually important region near seams (red circles). Because our method steers seams toward occluded regions, the checkerboard pattern appears substantially more continuous from typical camera viewpoints. By contrast, baselines exhibit visible seam artifacts in the zoomed-in insets.
Figure 9: Extended version of Fig. 4 in the main paper. Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara (Zhao et al., 2025), and OptCuts (Li et al., 2018) (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.
Figure 10: Extended version of Fig. 4 in the main paper. Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara (Zhao et al., 2025), and OptCuts (Li et al., 2018) (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.